OpenStudy (anonymous):

lim x->0 (x)/(|x|) Derivative of |x|?

OpenStudy (anonymous):

Hypothetically, looking at the graph of \(|x|\), then the derivative would be -1 in the interval \((-\infty,0)\) and 1 in the interval \((0,\infty)\) this is known as the signum function and is defined as: \[\sigma(x)=\frac{|x|}{x}\]

OpenStudy (anonymous):

Consider positive and negative values of \(x\). In other words, check the one-sided limits. For negative \(x\), you'd be approaching 0 from the left: \(\displaystyle\lim_{x\to0^-}\frac{|x|}{x}\). Likewise, for positive \(x\), you'd approach 0 from the right: \(\displaystyle\lim_{x\to0^+}\frac{|x|}{x}\). What do you know about \(|x|\) for negative and positive values of \(x\) ?

OpenStudy (anonymous):

So this limit does not exist?

OpenStudy (anonymous):

Sorry, I meant \(\dfrac{x}{|x|}\)

OpenStudy (anonymous):

because from the left it is -1 from the right it is 1?

OpenStudy (anonymous):

Right, the limit doesn't exist. The derivative is given by the sign function posted earlier.

OpenStudy (anonymous):

Thanks @SithsAndGiggles and @KeithAfasCalcLover